### v1

# Protected gap closing and reopening in topological-insulator Josephson junctions

Abstract
In the seminal proposal by Fu and Kane, the superconducting proximity effect was predicted to transform the surface state of a topological insulator (TI) into a topological superconductor, forming a nonchiral 1D Majorana state within a linear Josephson junction on the TI surface. The hallmark of this 1D Majorana state is a robust gap closing as a function of the superconducting phase difference $\varphi$ across the junction, which alternates in and out of the topological phase. These topological phase-transitions occur at $\varphi=(2n+1)\pi$ with integer $n$, leading to a $4\pi$-periodicity of the ground state. While the $4\pi$-periodicity has been indirectly inferred in the AC Josephson effect, the direct observation of the 1D Majorana state in a TI Josephson junction has remained contentious. Here, we report the direct observation of topological phase-transitions in a TI Josephson junction, where the local density of states is probed via tunnel contacts and $\varphi$ is controlled by a flux loop. The observed transitions are independent of the chemical potential, reinforcing their topological origin. Under an applied perpendicular magnetic field, Josephson vortices form, making $\varphi$ position-dependent. In this case, the gap closing occurs locally at the Josephson vortex cores where $\varphi=(2n+1)\pi$, which we also observe. Our findings provide direct confirmation of the Fu-Kane proposal and robust evidence for the emergence of topological superconductivity in a TI Josephson junction.
{affiliations}
Physics Institute II, University of Cologne, Zülpicher Str. 77, 50937 Köln, Germany
Institute for Quantum Information, RWTH Aachen University, 52056 Aachen, Germany
Max Planck Institute for the Science of Light, 91058 Erlangen, Germany
Department of Physics, Friedrich-Alexander-Universität Erlangen-Nürnberg, 91058 Erlangen, Germany
These authors contributed equally
1 Introduction
After more than 100 years from the discovery, superconductors remain a fascinating subject of research in physics. They become even more interesting when topology comes into play[1]. A special type of superconductor, called “spinless chiral $p$-wave superconductor”, is a prototypical topological superconductor hosting Majorana zero-modes at point-defects[2, 1]. The Majorana zero-modes obey non-Abelian exchange statistics[3, 4], which make them a promising candidate for topological quantum computing[5]. Due to this exciting prospect for applications, the realization of topological superconductivity and the observation of Majorana zero-modes have become major themes in condensed matter physics.
Since a spinless chiral $p$-wave superconductor is not yet found in nature, there have been significant efforts to effectively realize it in a hybrid system[6, 7]. Theoretically, a hybrid of a strongly spin-orbit-coupled material and a conventional superconductor would effectively realize the spinless chiral $p$-wave pairing due to spin-momentum locking. The earliest of such proposals made by Fu and Kane in 2008 is based on topological insulator (TI) surface states proximitised by a conventional $s$-wave superconductor[8]. However, even though the spin-momentum locking is most naturally realized in TIs, the experimental evidence for the realization of topological superconductivity obtained so far in this platform is weak and relies on indirect signatures, such as modified Fraunhofer patterns[9, 10, 11] and the microwave response[12, 13, 14] of a Josephson junction made on a TI, or a broad zero-bias conductance peak (ZBCP) observed in the Abrikosov vortex core of a proximitised TI[15]. The search for Majorana zero-modes has been more advanced on one-dimensional (1D) semiconducting Rashba nanowires, proximity-coupled to a superconductor, and exposed to a strong magnetic field[7]; however, despite a remarkable experimental progress[16, 17], the distinction between topological and trivial ZBCPs remains contentious in this 1D platform[18, 19, 20]. It has been realized early on[21] and emphasized in the recent literature[22, 23, 24, 25] that the direct observation of the topological phase-transition, the closing and reopening of the superconducting (SC) gap, is a more robust signature of a topological superconductor and thus Majorana zero-modes.
In this work, we turn to the TI platform and report the observation of protected gap closing and reopening in a TI Josephson junction upon changing the phase difference $\varphi$ across the junction, the key signature of the 1D Majorana state predicted by Fu and Kane for a phase-biased TI Josephson junction realizing 1D topological superconductivity[8]. The gap closing and reopening correspond to a switch in the ground-state parity, which implies a change in the $\mathbb{Z}_{2}$ topological invariant [2] and signals a topological phase-transition of the 1D Majorana state formed in the junction. The phase transition is topologically-protected in the absence of parity-mixing terms. At the phase transition, which occurs at $\varphi=(2n+1)\pi$ with integer $n$, the gap closes and a gapless 1D mode appears. Note that the Fu-Kane theory is only valid for a constant $\varphi$ along the junction. When a perpendicular magnetic field is applied to the junction, rendering $\varphi$ position-dependent[26], the gapless 1D mode becomes confined at the core of a Josephson vortex, where the local phase difference is $\varphi=(2n+1)\pi$. Because of this, the gapless 1D mode evolves into a Majorana zero-mode that is bound to the point defect formed by the Josephson vortex. In our experiment, we observe a periodic closing and reopening of the gap by tuning the phase difference across the junction in a SQUID geometry and simultaneously detecting the local density of states (LDOS) at different locations. The gap closing is accompanied by a multitude of low-lying excited states, as expected from theory. We further elucidate the locality of the involved states created by the phase gradient in perpendicular magnetic fields. Our observation is a direct experimental proof of the Fu-Kane theory and is a strong indication of the emergence of topological superconductivity in a TI Josephson junction. Based on this experiment, we discuss the desirable parameter range to realize well-protected Majorana zero-modes in a TI Josephson junction, which turns out to be within experimental reach.
2 Phase-biased topological-insulator Josephon junction
The 1D Majorana state predicted by Fu and Kane has a $4\pi$-periodic $\varphi$-dependence at zero momentum as shown in Fig. 1a. The corresponding evolution of the gap in the junction is depicted in Fig. 1c: The gap is trivial for $0\leq\varphi<\pi$. It closes at the topological phase-transition with $\varphi=\pi$ where the gapless 1D mode (green line) appears, before reopening in the topological regime, with $\pi<\varphi<3\pi$. In the topological regime the parity of the ground state is odd, meaning that the bound state below the chemical potential $\mu$ is filled with an unpaired electron (shown in yellow). Because of this, the junction returns to the initial state only at $\varphi=4\pi$. In contrast, for a conventional Andreev bound state, a $2\pi$-periodicity is expected, as depicted in Fig. 1b. Here, the dispersion is given by[27]
$E=\pm\Delta\sqrt{1-\tau\sin^{2}\left(\varphi/2\right)}\,\,,$ (1)
with $\tau$ the transmittance of the state and $\Delta$ the induced SC gap in the TI surface directly beneath the SC electrodes. In such a scenario, the gap can only close for the fine-tuned case of perfect transmission ($\tau$ = 1). Moreover, in this case, the ground-state parity would remain even across the gap closing as a conventional Andreev bound state is doubly degenerate due to the spin degeneracy.
A scanning electron microscope (SEM) image of our typical device is shown in Fig. 1d. A Josephson junction is created by a thin film of Al on top of an exfoliated flake of the bulk-insulating TI material BiSbTeSe2[28, 29]. The two sides of the junction are connected by an Al loop, which allows us to accurately tune the phase difference $\varphi$ by applying a perpendicular magnetic field. Such a device geometry is commonly used to study the bound state spectrum in an SNS-type Josephson junction via tunnel spectroscopy[30, 31, 32, 33]. Normal electrodes made of Au are deposited after putting a thin layer of $\mathrm{Al_{2}O_{3}}$ as a tunnel barrier. Due to the native oxide of Al, its surface can be treated as an insulator and the overlap to the Au electrodes does not contribute to the transport. A back-gate voltage can be applied through the degenerately-doped silicon substrate. Therefore, this setup enables us to directly probe the LDOS of the TI Josephson junction and to test the Fu-Kane theory[8]. Note that when the perpendicular magnetic field is very weak such that the large loop is threaded by only a few flux quanta (and no Josephson vortex nucleates in the junction), the phase $\varphi$ is almost constant along the junction, a situation akin to the Fu-Kane model. The case at larger magnetic field, when Josephson vortices nucleate in the junction, is discussed later. We focus on the results obtained from a device similar to the one shown in Fig. 1d with two independent tunnel junctions TJ1 and TJ2 located at different positions on the same Josephson junction. Additional data obtained from a second, essentially identical device, is shown in the supplement to demonstrate the reproducibility of our results.
3 Tunnel spectra and protected gap closing
In Fig. 1e, we show the continuous evolution of the differential conductance $dI/dV$ of TJ1 with the applied DC bias voltage $V_{b}$ and the perpendicular magnetic field $B$ around zero. Line cuts for selected fields are shown in Fig. 1f. Without magnetic field applied, we observe sharp particle-hole symmetric peaks at the bias voltages of ±$\pm 130\,\mathrm{\mu eV}$ and a strong reduction in $dI/dV$ at lower bias. The constant $dI/dV$ at higher bias corresponds to a resistance of about $400\,\mathrm{k\Omega}$, indicating we are in the tunnelling regime, where the observed $dI/dV$ is proportional to the LDOS. The peak energy of $130\,\mathrm{\mu eV}$ is smaller than the SC gap of our Al-film, which can be calculated from the critical temperature to be $170\,\mathrm{\mu eV}$. This confirms that the LDOS of the proximitised TI surface, not the LDOS of the Al electrode, is being measured. The LDOS at low bias does not drop to zero but remains finite, which indicates additional in-gap states that are not fully proximitised, consistent with our previous result on similar TI Josephson junctions[34]. Such a soft gap has also been reported for comparable devices based on carbon nanotube[30], graphene[31], HgTe[32], or InAs[33].
As a finite phase bias $\varphi$ is generated across the junction with the application of a perpendicular magnetic field, the positions of the particle-hole symmetric peaks shift to a lower bias voltage and their amplitude is diminished. For the magnetic field of about $115\,\mathrm{\mu T}$ that corresponds to $\varphi=\pi$, the particle-hole symmetric peaks merge into a single broad peak at zero bias, signalling the closing of the gap. The spectrum then evolves periodically, with the conductance at $230\,\mathrm{\mu T}$ becoming identical to the one observed at zero field. Having the Fu-Kane theory in mind, our data is consistent with the interpretation that the peak positions trace the gap, corresponding to the lowest-lying state at zero momentum, that evolves with $\varphi$ and closes at $\pi$ (see Fig. 1a). Above this gap (but still below the bulk gap), there is a continuum of states corresponding to the motion along the junction[8]. Our LDOS spectra reflects this physics showing a higher conductance above the gap when compared to the value below. At $\varphi=\pi$, the gap closing indicates a topological phase-transition[8].
It is important to note that the gap closing is observed despite the moderate interface transparency of our junction. Moderate transparency indicates that the junction transmittance $\tau$ is less than 1, which implies that conventional Andreev bound states should be gapped according to Eq. (1). Indeed, similar devices made on various materials reported in the literature[30, 31, 32, 33] all had $\tau<1$ and exhibited gapped spectra at $\varphi=\pi$. In contrast, our observation shows a gap closing, which is distinctly different from these results in non-topological materials supporting that the gap closing is of topological origin.
To further corroborate this interpretation, we have performed numerical simulations of the differential conductance $G$ expected from the 1D Majorana state for the geometry and parameters of our devices (see supplement for details). Figure 2a shows the calculated $G$ as a function of $V_{b}$ and $\varphi$, while the line cuts for the same $\varphi$ values as in Fig. 1f are shown in Fig. 2b. These simulations clearly identify the two branches that cross at $\varphi=(2n+1)\pi$ to be the even- and odd-parity bound states having a $4\pi$-periodicity. In Fig. 2b, the coherence peaks show up at the edge of the excitation gap and they diminish rapidly with increasing $\varphi$. Most importantly, the simulations show that the gap closing at $\varphi=\pi$ results in a broad hump rather than a sharp ZBCP. All these results are consistent with our experimental observation.
4 Locality of the gap closing in perpendicular magnetic fields
Next, we study the behaviour of our device in higher magnetic fields and compare the LDOS probed at TJ1 and TJ2. When a perpendicular magnetic field $B$ is applied to the junction, $\varphi$ becomes dependent on the position along the junction ($y$ direction). If $B$ is strong enough to nucleate Josephson vortices, the position where the local phase difference becomes $\varphi=(2n+1)\pi$ is identified as the Josephson vortex-core, and the gapless 1D mode in the Fu-Kane theory for global $\varphi=(2n+1)\pi$ turns into a Majorana zero-mode confined to the Josephson vortex-core[26]. The concept of this Majorana zero-mode generation is sketched in Fig. 3a. A spatially varying phase $\varphi$ along a TI Josephson junction caused by $B$ leads to a position-dependent gap function $m\propto\cos\frac{\varphi(y)}{2}$ as indicated by the colour code in the junction[26]. Associated with the sign change in the gap function, the gap closes locally at $\varphi(y)=(2n+1)\pi$ (white area), corresponding to the core of the Josephson vortex, and a Majorana zero-mode (green dots) nucleates that is confined along the junction by the phase gradient. We note that the topologically-protected nature of the gap closing is intuitively understandable in this situation: the region of the junction with $-\pi<\varphi(y)<\pi$ (red) has the even-parity ground state with $\mathbb{Z}_{2}$ index of 0, whereas the region with $\pi<\varphi(y)<3\pi$ (blue) has the odd-parity ground state with $\mathbb{Z}_{2}$ index of 1; therefore, the boundary between the two regions must become gapless for the topological invariant to change.
We indeed observed the locality of this gap closing. The top and bottom panels of Fig. 3c show the $dI/dV$ data simultaneously measured at TJ1 and TJ2 as a function of $V_{b}$ and $B$ in the 5 – 6 mT range. Both junctions present a comparable evolution of their bound state spectrum, but there is a clear phase shift and their periods are slightly different. For TJ2 we find a periodicity of $200\,\mathrm{\mu T}$, smaller than that of TJ1 ($230\,\mathrm{\mu T}$). This difference in the period is consistent with the sample geometry (see supplement for details) in which the effective junction area considering the flux focusing effect[34, 35, 36] is about $\sim$10of the loop area. To understand the LDOS at TJ1 and TJ2, we have calculated the expected texture of the gap function based on the device geometry. In our junctions, the self-inductance is negligible and its width is much shorter than the Josephson penetration depth (see supplement), justifying the description of the phase gradient in terms of
$\varphi(y)=\frac{2\pi B}{\Phi_{0}}\left(A+yL_{\rm eff}\right)\,\,,$ (2)
with $A$ the loop area, $\Phi_{0}=\frac{h}{2e}$ the magnetic flux quantum, and $L_{\rm eff}$ the effective length of the junction taking the flux-focusing effect into account. The evolution of the gap function $m\propto\cos\frac{\varphi(y)}{2}$ is calculated from this formula with the loop and junction area as fitting parameters (see supplement) and visualized in the middle panel of Fig. 3c. We show snapshots of the phase-texture in the junction for a range of magnetic fields corresponding to the experimental situation. In this range, there are two to three Josephson vortices inside the junction, whose position is tuned by the global perpendicular magnetic field. They enter the junction from the bottom where TJ2 is located, and move upwards to TJ1 with increasing magnetic field. Whenever a Josephson vortex-core crosses the location of a tunnel junction, a gap closing is observed. This demonstrates the local nature of the gap closing in the presence of Josephson vortices. The data shown in Fig. 3c was taken at an applied back-gate voltage of $V_{\rm G}=-5$ V, and we chose the magnetic field range of 5 – 6 mT for the visibility of the phase gradient. The same pattern was found to continue into higher magnetic fields until the Al superconductivity was suppressed at 15 mT (see supplement).
We note that the spectra shown in Fig. 3c essentially look the same as those in Fig. 1e. This is because the Majorana zero-mode in the Josephson vortex-core is not well-separated in energy. For the conditions realized, it is accompanied by a quasi-continuum of low-energy bound states[26]. Taking the smearing due to the finite-temperature into account, the local spectrum becomes a broad peak, indistinguishable from the Fu-Kane situation. In Fig. 3b, we show the magnitude of the gap function near a Josephson vortex-core and the finite-energy bound states calculated from the model in Ref. 26, along with the spatial extent of the expected Majorana zero-mode. For the parameters relevant to our device with $\mu$ = 25 meV, the calculated finite-energy bound states form a quasi-continuum with a minigap of only 2% of $\Delta$ (i.e. 3 $\mu$eV), which is of the order of the thermal energy at the base temperature.
5 Stability of the gap closing
In the next step, we check the robustness of the observed features with respect to the chemical potential, as tuned by a global back gate. Figure 4a shows the $dI/dV$ data in $B$ = 0 T at a high bias voltage ($V_{\rm b}$ = 1 mV) as a function of the back-gate voltage $V_{G}$ for both TJ1 and TJ2. The overall tendency is that the conductance decreases for more negative gate-voltages, indicating that the TI surface is in the $n$-type regime. Superimposed to this general trend, we observe broad peaks (typical peak width of a few V), which are commonly found in dielectric-based tunnelling devices and conventionally explained by resonant states in the dielectric[31]. Figure 4b presents the $V_{\rm G}$-dependencies of the normalized conductance spectra obtained at TJ1: the upper and lower panels show the evolutions of the maximum excitation gap for $\varphi=0$ (upper) and the spectrum observed for $\varphi=\pi$ (lower) in a wide range of $V_{\rm G}$ from $-20$ to $+20$ V. The maximum excitation gap is very stable, indicating that the resonance states due to impurities are not affecting the proximity-induced gap. On the other hand, the position of the conductance peak for $\varphi=\pi$ is slightly shifted to the negative $V_{\rm b}$ side at $V_{\rm G}\approx$ 12 – 20 V, which is probably due to an influence of the strong resonance states located around 13 V (see Fig. 4a). Remarkably, the broad ZBCP that accompanies the gap closing is stable for a wide range of $V_{\rm G}$ from $-10$ V to $+12$ V. It is worth noting that the ZBCP observed in this $V_{\rm G}$ range is not correlated with the weaker resonance states seen in Fig. 4a, which further supports the intrinsic origin of the gap closing. For $V_{\rm G}\lesssim-10$ V, the gap closing is not accompanied by a clear ZBCP. We believe that this is because our non-tunable tunnel contacts did not couple sufficiently well to the bound states in this regime, preventing us from resolving a peak.
In Figs. 4c and 4d, we show the $B$-dependencies of the normalized conductance spectra for equally-spaced $V_{\rm G}$ values between $-20$ and $+20$ V for both TJ1 and TJ2. One can immediately see that the basic magnetic-field-periodic change in $dI/dV$ is very stable at all $V_{\rm G}$ values. However, upon close inspection, the occasional breaking of exact electron-hole symmetry is visible. Particularly in TJ1, the tunnel-coupling of the electrode to the electron and hole states in the TI surface is modified by back gating: At $V_{\rm G}=-10\,\mathrm{V}$, the electron-like bound states give a stronger signal than the hole-like ones, whereas the situation is reversed for $V_{\rm G}=+15\,\mathrm{V}$. Similar behaviour has been reported for comparable devices, e.g., based on carbon nanotubes[30], and it was explained to be due to state-dependent tunnel couplings. Overall, we conclude that the gap closing is robust with respect to changes of the chemical potential indicating that we observe a topological phase-transition.
In passing, we mention that in a device where the tunnel barrier was not in a favourable condition to probe the TI surface, we observed strongly dispersing bound states in a very narrow range of $V_{\rm G}$ associated with a spurious ZBCP occurring at a particular $V_{\rm G}$ value (see supplement). These features are consistent with the Andreev quantum-dot states predicted in Ref. 37, which might occur in the dielectric. The observations of two different types of ZBCPs, stable and spurious, give confidence that the former is associated with an intrinsic gap closing.
6 Discussions
Our observation of a stable gap closing and reopening in a TI Josephson junction having only a moderate transparency cannot be understood in terms of conventional Andreev bound states and gives the first direct proof of the topological phase-transition predicted by the Fu-Kane theory. To firm up the appearance of topological superconductivity, it would be an interesting future experiment to employ a charge sensing technique[38] in order to detect the parity change in the junction before and after the gap closing. Of course, the definitive proof of topological superconductivity is the observation of Majorana zero-modes, well-protected by a substantial excitation gap, and the demonstration of their non-Abelian exchange statistics through braiding. In a TI Josephson junction, well-protected Majorana zero-modes can in principle be generated in Josephson vortex-cores as we discuss below. Once such protected Majorana zero-modes are generated, they can be physically moved along the junction in a controlled manner[39]; this degree of freedom allows us to conceive braiding operations in a three-slit Josephson junction[39] and the resulting change in the parity can be detected by charge sensing[38]. Another possibility to confirm the non-Abelian statistics is to build a Josephson-vortex interferometer proposed by Grosfeld and Stern[40], which detects the result of braiding using the Aharonov-Casher effect.
Let us estimate the necessary parameter range to obtain well-protected Majorana zero-modes in Josephson vortex-cores. There are two quantities that are crucial for this consideration: One is the spatial extent of the Majorana zero-mode $\lambda_{\rm B}=\sqrt{v_{\rm M}l_{B}/\pi\Delta}$, where $l_{B}=\Phi_{0}/(BL_{\rm eff})$ is the distance between Josephson vortices in the junction and $v_{\rm M}$ is the renormalized Majorana velocity. To calculate $v_{\rm M}$, we use the formula (see supplement for details)
$v_{\rm M}=v_{\rm F}\left(\frac{\Delta^{2}}{\Delta^{2}+\mu^{2}}\right)\frac{\cos\left(\frac{\mu L}{\hbar v_{\rm F}}\right)+\frac{\Delta}{\mu}\sin\left(\frac{\mu L}{\hbar v_{\rm F}}\right)}{1+\frac{\Delta L}{\hbar v_{\rm F}}}$ (3)
valid in the short-junction regime $L\lesssim\hbar v_{\rm F}/\Delta$ ($L$ is the true length of the junction). For large $\mu\gg\Delta$, the formula goes over to the simple expression $v_{\rm M}\simeq v_{\rm F}(\Delta/\mu)^{2}$ given in Ref. 26. The other important quantity is the energy of the first excited state, which is given as $E_{1}\approx\sqrt{2\pi v_{\rm M}\Delta/l_{B}}$ near $\varphi=(2n+1)\pi$ [26]. Note that reducing $\mu$ enhances $v_{\rm M}$, and thus $E_{1}$ and $\lambda_{\rm B}$. However, $\lambda_{\rm B}$ must be shorter than $l_{B}/2$ to avoid the overlap of neighbouring Majorana zero-modes which bounds $\mu$ from below. To see this competition, we theoretically consider the case of using Nb for the SC electrodes. We calculate the dependencies of $\lambda_{\rm B}$ and $E_{1}$ on $\mu$ and $B$. The results, shown in Figs. 2c and 2d, indicate that a moderately large $\mu$ is indeed helpful to realize a short $\lambda_{\rm B}$ and the magnetic field helps to obtain a usefully large $E_{1}$. In Fig. 2d, we draw the boundaries to guarantee a sufficient separation of Majorana zero-modes ($\lambda_{\rm B}/l_{B}<0.2$, blue line) and a sufficiently large excitation gap ($E_{1}/\Delta>0.2$, green line). In the region between the blue and green lines, see Fig. 2d, Majorana zero-modes are well-protected. In fact, these estimates suggest that it is not necessary to tune $\mu$ very close to the Dirac point as long as a magnetic field of the order of 100 mT is used to generate Josephson vortices — an encouraging result for the prospects of Majorana zero-modes in this platform. It is prudent to note that the soft gap, which we observed in our devices, is a problem for braiding, and a hard gap must be realized by improving the interface transparency between the TI and a superconductor. In this regard, we envision the use of the self-epitaxy technique discovered for proximitising (Bi1-xSbx)2Te3 thin films with the PdTe2 superconductor, because a very high transparency of 0.96 was achieved with this technique[41].
In conclusion, we have observed periodic gap-closing and reopening in a phase-biased TI Josephson junction via tunnelling spectroscopy. We have shown that this phenomenon is robust as a function of an extended parameter range in the chemical potential and the perpendicular magnetic field. The robustness of the observed gap closing and reopening is in sharp contrast to the signal expected for a conventional moderate-transparency Josephson junction or other types of zero-energy states which originate from Andreev quantum-dot states[37]. The robustness of the gap closing and reopening is a key feature of the topological phase-transition that occurs in the 1D Majorana state as a function of $\varphi$. In the magnetic-field range where a couple of Josephson vortices nucleate in the junction, we have observed the locality of the gap closing implying a confinement of the zero-energy mode[26]. However, within the magnetic-field range where Al remains superconducting, an abundance of low-lying excited states arises and this leads to spectra that are indistinguishable from the gap closing of an extended 1D Majorana mode. We have discussed the necessary parameter range to realize well-protected Majorana zero-modes in the Josephson vortex core. Intriguingly, we found that maintaining a chemical potential $\mu$ not too close to the Dirac point is crucial to avoid Majorana delocalisation; a $\mu$ of approximately 25 meV facilitates the realization of well-protected Majorana zero-modes in a magnetic field of 100 mT accessible with a Nb electrode. Considering the high controllability of the Josephson vortex, this result is promising for braiding experiments. Overall, our experiment elucidates the topological nature of a TI Josephson junction and underscores the realistic prospect of generating well-protected Majorana zero-modes in this system.
{methods}
Materials and device fabrications: Bulk single crystals of BiSbTeSe2 were synthesized by the modified Bridgman method using high-purity (99.9999%) Bi, Sb, Te, and Se as starting materials by following the recipe described in Ref. 28. We mechanically exfoliated ultra-thin BiSbTeSe2 flakes from bulk single crystals on degenerately-doped Si wafers covered by 280 nm SiO2. The latter acts as dielectric for the back gate. BiSbTeSe2 flakes suitable for device fabrication were identified under an optical microscope. The junctions were defined by electron beam lithography. PMMA A4 resist was exposed using a Raith PIONEER Two system. After development, residual resist in the contact areas was removed by a gentle oxygen plasma treatment. Subsequent cleaning in a hydrochloric acid solution shortly before metallisation ensured the removal of any oxide layer. Then, 3 nm + 45 nm Ti/Al contacts were thermally evaporated after a gentle in-situ Ar cleaning and the device was exposed to air to form a native Al2O3 layer. An additional atomic layer deposition of a thin Al2O3 layer onto the whole device was performed at 80 C using Ultratec Savannah S200 to form a tunnel barrier. For the fabrication of the tunnel electrodes, another electron beam lithography was performed to define the pattern. A 5-nm-thick Pt layer was sputter-deposited to act as wetting layer, followed by sputtering an additional 45-nm-thick Au layer. The contact resistances achieved by this procedure are typically larger than $100\mathrm{k\Omega}$, ensuring electron tunnelling to be the prevalent form of transport. The precise device geometry was determined by scanning electron microscopy after the measurements were finished.
Measurements: Transport measurements were performed at base temperature ($\sim$50 mK) of our dry dilution refrigerator (Oxford Instruments TRITON 300). Electrical lines incorporate RC and copper-powder filters to reduce noise. Using a two-probe configuration, the differential conductance $dI/dV$ was measured between each individual tunnel electrode and the SC electrode with a standard low-frequency lock-in technique using a Basel instruments current preamplifier. A small DC-offset introduced by the preamplifier was numerically corrected for after the measurements. An AC voltage of 5–10 $\mu$V was superimposed to a DC bias voltage. For investigations of the magnetic field dependence, we used our integrated 6/1/1-T superconducting vector magnet.
Simulations: To qualitatively reproduce the conductance spectrum of the TI Josephson junction, a transport simulation was performed using the Python library Kwant[42]. As a model for the three-dimensional strong TI, the Bernevig-Hughes-Zhang model was chosen and implemented on a cubic lattice. Superconductivity was included in the Bogoliubov–de-Gennes formalism. The system has been extended infinitely along the junction and probed by a tunnel lead attached perpendicularly to the junction. For more information see supplement.
{addendum}
This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant Agreement No. 741121) and was also funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1 - 390534769, as well as by the DFG under CRC 1238 - 277146847 (Subprojects A04 and B01).
MB grew the BSTS2 crystals, JS and ENN fabricated the devices, performed the experiments and analysed the data together with YA. AM and AZ performed numerical calculations, FH and YA conceived and supervised the project. JS, FH, and YA wrote the manuscript with input from all authors.
The authors declare no competing interests.
Correspondence and requests for materials should be addressed to Y.A. (ando@ph2.uni-koeln.de).
Data availability: The data that support the findings of this study are available at the online depository zenodo with the identifier 10.5281/zenodo.11615514 and in Supplementary Information.

### v2

@@
-Protected gap closing and reopening in topological-insulator Josephson junctions
+thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work
 Abstract
-In the seminal proposal by Fu and Kane, the superconducting proximity effect was predicted to transform the surface state of a topological insulator (TI) into a topological superconductor, forming a nonchiral 1D Majorana state within a linear Josephson junction on the TI surface.
-The hallmark of this 1D Majorana state is a robust gap closing as a function of the superconducting phase difference $\varphi$ across the junction, which alternates in and out of the topological phase.
-These topological phase-transitions occur at $\varphi=(2n+1)\pi$ with integer $n$, leading to a $4\pi$-periodicity of the ground state.
-While the $4\pi$-periodicity has been indirectly inferred in the AC Josephson effect, the direct observation of the 1D Majorana state in a TI Josephson junction has remained contentious.
-Here, we report the direct observation of topological phase-transitions in a TI Josephson junction, where the local density of states is probed via tunnel contacts and $\varphi$ is controlled by a flux loop.
-The observed transitions are independent of the chemical potential, reinforcing their topological origin.
+In the seminal proposal by Fu and Kane, the superconducting proximity effect is used to realize topological superconductivity in the topological surface state (TSS) of a 3D topological insulator (TI).
+In a line Josephson junction made on the TI surface, the spin-momentum locking of the TSS guarantees the existence of a pair of spin-non-degenerate, perfectly transmitted Andreev modes.
+These modes lead to robust gap closing and parity alteration as a function of the superconducting phase difference $\varphi$ across the junction.
+Here, we report the observation of the predicted gap closing at $\varphi=(2n+1)\pi$ in a TI Josephson junction ($n$ integer), where the local density of states is probed via tunnel contacts and $\varphi$ is controlled by a flux loop.
+This phenomenon is robust for a wide range of chemical potentials, supporting its TSS origin.
 Under an applied perpendicular magnetic field, Josephson vortices form, making $\varphi$ position-dependent.
 In this case, the gap closing occurs locally at the Josephson vortex cores where $\varphi=(2n+1)\pi$, which we also observe.
-Our findings provide direct confirmation of the Fu-Kane proposal and robust evidence for the emergence of topological superconductivity in a TI Josephson junction.
-{affiliations}
-Physics Institute II, University of Cologne, Zülpicher Str. 77, 50937 Köln, Germany
-Institute for Quantum Information, RWTH Aachen University, 52056 Aachen, Germany
-Max Planck Institute for the Science of Light, 91058 Erlangen, Germany
-Department of Physics, Friedrich-Alexander-Universität Erlangen-Nürnberg, 91058 Erlangen, Germany
-These authors contributed equally
-1 Introduction
-After more than 100 years from the discovery, superconductors remain a fascinating subject of research in physics.
+Our results confirm the fundamental role of spin-momentum locking in the Andreev physics in the TSS, which implies that the observed gap closing and reopening has a topological nature.
+I Introduction
+After more than 100 years from the discovery, superconductors remain a fascinating subject in physics.
 They become even more interesting when topology comes into play [1].
@@
 Theoretically, a hybrid of a strongly spin-orbit-coupled material and a conventional superconductor would effectively realize the spinless chiral $p$-wave pairing due to spin-momentum locking.
-The earliest of such proposals made by Fu and Kane in 2008 is based on topological insulator (TI) surface states proximitised by a conventional $s$-wave superconductor[8].
-However, even though the spin-momentum locking is most naturally realized in TIs, the experimental evidence for the realization of topological superconductivity obtained so far in this platform is weak and relies on indirect signatures, such as modified Fraunhofer patterns[9, 10, 11] and the microwave response[12, 13, 14] of a Josephson junction made on a TI, or a broad zero-bias conductance peak (ZBCP) observed in the Abrikosov vortex core of a proximitised TI[15].
+The earliest of such proposals made by Fu and Kane in 2008 is based on the topological surface state (TSS) of a 3D topological insulator (TI) proximitised by a conventional $s$-wave superconductor [8].
+However, even though the spin-momentum locking is most naturally realized in TIs, the experimental evidence for topological superconductivity obtained so far in this platform is weak and relies on indirect signatures, such as modified Fraunhofer patterns [9, 10, 11] and the microwave response [12, 13, 14] of a Josephson junction made on a TI, or a broad zero-bias conductance peak (ZBCP) observed in the Abrikosov vortex core of a proximitised TI [15].
 The search for Majorana zero-modes has been more advanced on one-dimensional (1D) semiconducting Rashba nanowires, proximity-coupled to a superconductor, and exposed to a strong magnetic field [7]; however, despite a remarkable experimental progress [16, 17], the distinction between topological and trivial ZBCPs remains contentious in this 1D platform [18, 19, 20].
-It has been realized early on[21] and emphasized in the recent literature[22, 23, 24, 25] that the direct observation of the topological phase-transition, the closing and reopening of the superconducting (SC) gap, is a more robust signature of a topological superconductor and thus Majorana zero-modes.
-In this work, we turn to the TI platform and report the observation of protected gap closing and reopening in a TI Josephson junction upon changing the phase difference $\varphi$ across the junction, the key signature of the 1D Majorana state predicted by Fu and Kane for a phase-biased TI Josephson junction realizing 1D topological superconductivity[8].
-The gap closing and reopening correspond to a switch in the ground-state parity, which implies a change in the $\mathbb{Z}_{2}$ topological invariant [2] and signals a topological phase-transition of the 1D Majorana state formed in the junction.
-The phase transition is topologically-protected in the absence of parity-mixing terms.
-At the phase transition, which occurs at $\varphi=(2n+1)\pi$ with integer $n$, the gap closes and a gapless 1D mode appears.
+It was recognized early on [21] and emphasized in the recent literature [22, 23, 24, 25] that a direct observation of the closing and reopening of the superconducting (SC) gap, which implies the topological phase-transition, is a more robust signature of topological superconductivity and thus Majorana zero-modes.
+Figure 1: Phase-dependent Andreev bound states in a TI Josephson junction.
+a,b, Phase-dispersions of the perpendicular ($k_{y}$ = 0) mode in a TI line junction (a) and conventional Andreev bound states (ABS) with $\tau=0.9$ (b).
+In panel a, the red and blue lines correspond to the even- and odd-parity bound states, respectively.
+c, Evolution of the topological phase in the junction; here, the junction area is colored according to the gap function $m\propto\cos\frac{\varphi(y)}{2}$ (see the color scale).
+As $\varphi$ is tuned, the nature of the gap changes from trivial ($\varphi$ = 0, red) to topological ($\varphi$ = $2\pi$, blue).
+This is accompanied by a change in ground state parity from even to odd, as indicated by the extra electron (yellow) in the junction.
+In-between, the gap has to close ($\varphi=\pi$) via a gapless Majorana mode (green line).
+As the phase difference reaches $\varphi=4\pi$, a full period is completed.
+d, False-color SEM image of a typical device.
+The exfoliated TI flake (red) is contacted by Al (blue) forming an RF-SQUID geometry.
+Metallic contacts (Au) allow for probing the LDOS at two positions (TJ1 and TJ2) in the junction.
+e, color mapping of $dI/dV$ in the $B$ vs $V_{\rm b}$ plane measured on TJ1 at 50 mK with $V_{\rm G}$ = 0 V.
+At $\varphi$ = 0 (mod $2\pi$), a proximity-induced gap opens with conductance peaks at $\pm$130 $\mu$V.
+The gap closes and a broad ZBCP is observed at $\varphi=\pi$.
+A periodicity of 230 $\mu$T is observed, consistent with the device geometry.
+f, Line-cuts of the data in e at three selected magnetic fields corresponding to $\varphi$ = 0, $\frac{3}{4}\pi$, and $\pi$.
+In this work, we turn to the TI platform and report the observation of robust gap closing and reopening in a TI Josephson junction upon changing the phase difference $\varphi$ across the junction.
+The spin-momentum locking of the TSS prohibits exact back-scattering, and hence the normal reflection is forbidden for a perpendicular mode at the boundary to a superconductor.
+Such an absence of normal back-scattering in turn guarantees perfect Andreev reflection, leading to a pair of perfectly transmitted, non-degenerate Andreev bound states (ABSs) in a short Josephson junction.
+These perfectly transmitted modes become gapless for $\varphi=\pi$, and the spin-non-degeneracy of the TSS makes the electron and hole branches to have opposite parity [26].
+As $\varphi$ is varied, these nondegenerate modes cross and the ground-state parity alternates at $\varphi=(2n+1)\pi$ with $n$ integer [8].
+Since the fermion parity is the $\mathbb{Z}_{2}$ topological invariant of a 1D superconductor [2], the gap closing and reopening corresponds to a topological phase transition.
+It is important to emphasize that the appearance of the perfectly transmitted ABSs is a direct consequence of the spin-momentum locking of the TSS that prohibits back-scattering.
+Therefore, as long as the Josephson effect takes place in the TSS, the topological phase transition as a function of $\varphi$ is a necessity.
+Nevertheless, this fundamental property has not been experimentally confirmed so far in TI Josephson junctions.
 Note that the Fu-Kane theory is only valid for a constant $\varphi$ along the junction.
-When a perpendicular magnetic field is applied to the junction, rendering $\varphi$ position-dependent[26], the gapless 1D mode becomes confined at the core of a Josephson vortex, where the local phase difference is $\varphi=(2n+1)\pi$.
+When an out-of-plane magnetic field is applied to the junction, rendering $\varphi$ position-dependent [27], the gapless 1D mode becomes confined at the core of a Josephson vortex, where the local phase difference is $\varphi=(2n+1)\pi$.
 Because of this, the gapless 1D mode evolves into a Majorana zero-mode that is bound to the point defect formed by the Josephson vortex.
-In our experiment, we observe a periodic closing and reopening of the gap by tuning the phase difference across the junction in a SQUID geometry and simultaneously detecting the local density of states (LDOS) at different locations.
-The gap closing is accompanied by a multitude of low-lying excited states, as expected from theory.
+In our experiment, we observe periodic closings and reopenings of the gap by tuning $\varphi$ across the junction in a SQUID geometry and simultaneously detecting the local density of states (LDOS) at different locations.
+The gap closing, which is confirmed to be robust against a wide range of chemical potential, is accompanied by a multitude of low-lying excited states, as expected from theory.
 We further elucidate the locality of the involved states created by the phase gradient in perpendicular magnetic fields.
-Our observation is a direct experimental proof of the Fu-Kane theory and is a strong indication of the emergence of topological superconductivity in a TI Josephson junction.
-Based on this experiment, we discuss the desirable parameter range to realize well-protected Majorana zero-modes in a TI Josephson junction, which turns out to be within experimental reach.
-2 Phase-biased topological-insulator Josephon junction
-The 1D Majorana state predicted by Fu and Kane has a $4\pi$-periodic $\varphi$-dependence at zero momentum as shown in Fig.
-1a.
+Hence, our observations confirm the consequence of the spin-momentum locking in the TSS in a most direct manner.
+Based on this promising result, we suggest future experiments to nail down the topological superconductivity and to generate well-protected Majorana zero-modes.
+Figure 2: Theoretical simulations.
+a, Color-mapping of the simulated differential conductance $G$ in units of the conductance quantum $G_{0}=e^{2}/h$ as a function of $V_{b}$ and $\varphi$ for a temperature $T=\Delta/(30k_{\rm B})$ in the Fu-Kane regime with $\varphi$ constant along the junction.
+b, Simulated behavior of the differential conductance $G(V_{b})$ at $\varphi$ = 0, $\frac{3}{4}\pi$, and $\pi$, which correspond to those shown in Fig.
+1f.
+Note that in the simulation only the surface states and a single transversal mode are taken into account which explains the softer-gap in the experimental results of Fig.
+1.
+II Phase-biased TI Josephon junction
+The perpendicular mode of the ABSs (i.e.
+zero longitudinal momentum $k_{y}$) appearing in a line junction made on the TSS is expected to present a $4\pi$-periodic $\varphi$-dependence as shown in Fig.
+1a in the short-junction limit.
+As mentioned above, this is dictated by the spin-momentum locking which prohibits back-scattering.
 The corresponding evolution of the gap in the junction is depicted in Fig.
@@
 $E=\pm\Delta\sqrt{1-\tau\sin^{2}\left(\varphi/2\right)}\,\,,$ (1)
-with $\tau$ the transmittance of the state and $\Delta$ the induced SC gap in the TI surface directly beneath the SC electrodes.
-In such a scenario, the gap can only close for the fine-tuned case of perfect transmission ($\tau$ = 1).
-Moreover, in this case, the ground-state parity would remain even across the gap closing as a conventional Andreev bound state is doubly degenerate due to the spin degeneracy.
-A scanning electron microscope (SEM) image of our typical device is shown in Fig.
-1d.
-A Josephson junction is created by a thin film of Al on top of an exfoliated flake of the bulk-insulating TI material BiSbTeSe2[28, 29].
-The two sides of the junction are connected by an Al loop, which allows us to accurately tune the phase difference $\varphi$ by applying a perpendicular magnetic field.
-Such a device geometry is commonly used to study the bound state spectrum in an SNS-type Josephson junction via tunnel spectroscopy[30, 31, 32, 33].
+with $\tau$ the transmittance of the state and $\Delta$ the induced SC gap in the normal metal beneath the SC electrodes.
+In such a scenario, the gap can close in the case of perfect transmission ($\tau$ = 1), but the ground-state parity remains even across the gap closing, as conventional ABSs are doubly degenerate.
+[…]
 Normal electrodes made of Au are deposited after putting a thin layer of $\mathrm{Al_{2}O_{3}}$ as a tunnel barrier.
 Due to the native oxide of Al, its surface can be treated as an insulator and the overlap to the Au electrodes does not contribute to the transport.
-A back-gate voltage can be applied through the degenerately-doped silicon substrate.
+A back-gate voltage can be applied via the degenerately-doped silicon substrate.
 Therefore, this setup enables us to directly probe the LDOS of the TI Josephson junction and to test the Fu-Kane theory [8].
@@
 Additional data obtained from a second, essentially identical device, is shown in the supplement to demonstrate the reproducibility of our results.
-3 Tunnel spectra and protected gap closing
+III Tunnel spectra and gap closing
 In Fig.
@@
 This confirms that the LDOS of the proximitised TI surface, not the LDOS of the Al electrode, is being measured.
-The LDOS at low bias does not drop to zero but remains finite, which indicates additional in-gap states that are not fully proximitised, consistent with our previous result on similar TI Josephson junctions[34].
-Such a soft gap has also been reported for comparable devices based on carbon nanotube[30], graphene[31], HgTe[32], or InAs[33].
-As a finite phase bias $\varphi$ is generated across the junction with the application of a perpendicular magnetic field, the positions of the particle-hole symmetric peaks shift to a lower bias voltage and their amplitude is diminished.
+The LDOS at low bias does not drop to zero but remains finite, which indicates additional in-gap states that are not fully proximitised, consistent with our previous result on similar TI Josephson junctions, where the average transparency was $\tau_{\rm av}\approx 0.8$ [37].
+Such a soft gap has also been reported for comparable devices based on graphene [31], carbon nanotube [33], HgTe [34], or InAs [35].
+As a finite phase bias $\varphi$ is generated across the junction with the application of an out-of-plane magnetic field, the positions of the particle-hole symmetric peaks shift to a lower bias voltage and their amplitude is diminished.
 For the magnetic field of about $115\,\mathrm{\mu T}$ that corresponds to $\varphi=\pi$, the particle-hole symmetric peaks merge into a single broad peak at zero bias, signalling the closing of the gap.
@@
 1a).
-Above this gap (but still below the bulk gap), there is a continuum of states corresponding to the motion along the junction[8].
-Our LDOS spectra reflects this physics showing a higher conductance above the gap when compared to the value below.
-At $\varphi=\pi$, the gap closing indicates a topological phase-transition[8].
-It is important to note that the gap closing is observed despite the moderate interface transparency of our junction.
-Moderate transparency indicates that the junction transmittance $\tau$ is less than 1, which implies that conventional Andreev bound states should be gapped according to Eq.
+Above this gap (but still below the bulk gap), there is a continuum of states corresponding to finite momentum $k_{y}$ along the junction [8], which is indeed reflected in our LDOS spectra.
+Within this interpretation, the gap closing at $\varphi=\pi$ corresponds to a topological phase-transition [8].
+It is useful to note that similar devices made on various materials in the ballistic regime [33, 34, 35] with $\tau<1$ exhibited gapped spectra at $\varphi=\pi$, consistent with Eq.
 (1).
-Indeed, similar devices made on various materials reported in the literature[30, 31, 32, 33] all had $\tau<1$ and exhibited gapped spectra at $\varphi=\pi$.
-In contrast, our observation shows a gap closing, which is distinctly different from these results in non-topological materials supporting that the gap closing is of topological origin.
-To further corroborate this interpretation, we have performed numerical simulations of the differential conductance $G$ expected from the 1D Majorana state for the geometry and parameters of our devices (see supplement for details).
+The gap closing observed here is most likely due to the unique property of the TSS, which necessitates the existence of a $\tau=1$ mode.
+To further corroborate this interpretation, we have performed numerical simulations of the differential conductance $G$ expected for the parameters of our devices (see supplement for details).
 Figure 2a shows the calculated $G$ as a function of $V_{b}$ and $\varphi$, while the line cuts for the same $\varphi$ values as in Fig.
@@
 2b.
-These simulations clearly identify the two branches that cross at $\varphi=(2n+1)\pi$ to be the even- and odd-parity bound states having a $4\pi$-periodicity.
+These simulations confirm that the even- and odd-parity bound states having a $4\pi$-periodicity produce the two branches that cross at $\varphi=(2n+1)\pi$, as observed in experiments.
 In Fig.
 2b, the coherence peaks show up at the edge of the excitation gap and they diminish rapidly with increasing $\varphi$.
-Most importantly, the simulations show that the gap closing at $\varphi=\pi$ results in a broad hump rather than a sharp ZBCP.
-All these results are consistent with our experimental observation.
-4 Locality of the gap closing in perpendicular magnetic fields
-Next, we study the behaviour of our device in higher magnetic fields and compare the LDOS probed at TJ1 and TJ2.
+Our simulations suggest that the exact shape of the spectrum at $\varphi=\pi$ depends sensitively on the experimental conditions, and the gap closing can result in a nearly flat $dI/dV$ (see supplement).
+All these results are consistent with our experimental observations.
+It should be emphasized that the gap closing always occurs at $\varphi=(2n+1)\pi$, but it does not need to be accompanied by a ZBCP, according to our simulations.
+IV Locality of the gap closing in perpendicular magnetic fields
+Figure 3: Locality of the gap closing.
+a, Schematic drawing of a planar TI Josephson junction in a perpendicular magnetic field, which generates a position-dependent phase $\varphi(y)$ across the junction.
+The distribution of the topological gap function $m\propto\cos\frac{\varphi(y)}{2}$ is depicted in the junction with a color scale.
+The position of $\varphi(y)=(2n+1)\pi$ corresponds to the Josephson vortex-core, where the gap closes (white color) and a Majorana zero-mode (green dot) appears.
+b, Position-dependent gap $|m|=|\cos\frac{\varphi(y)}{2}|$ near a Josephson vortex-core in a magnetic field of $B$ = 5 mT, plotted with the excited states calculated from the model of Ref.
+27 for a chemical potential $\mu$ = 25 meV and $\Delta$ = 0.17 meV (thin horizontal lines), which form a quasi-continuum with a minigap of only 2% of $\Delta$.
+The vertical red lines mark the spatial extent of the expected Majorana zero-mode appearing around $y$ = 0.
+c, color mapping of $dI/dV$ in the $B$ vs $V_{\rm b}$ plane for TJ1 (top panel) and TJ2 (bottom panel) measured in the magnetic field range of 5 – 6 mT at 50 mK with $V_{\rm G}=-5$ V.
+Note the slight difference in the periodicity, which is 230 $\mu$T for TJ1 and 200 $\mu$T for TJ2.
+The middle panel depicts how the distribution of the topological gap function in the junction calculated with Eq.
+(2) evolves with magnetic fields; the position where the gap function locally becomes zero with $\varphi(y)=(2n+1)\pi$ corresponds to local gap closing (white regions).
+Next, we study the behavior of our device in higher magnetic fields and compare the LDOS probed at TJ1 and TJ2.
 When a perpendicular magnetic field $B$ is applied to the junction, $\varphi$ becomes dependent on the position along the junction ($y$ direction).
-If $B$ is strong enough to nucleate Josephson vortices, the position where the local phase difference becomes $\varphi=(2n+1)\pi$ is identified as the Josephson vortex-core, and the gapless 1D mode in the Fu-Kane theory for global $\varphi=(2n+1)\pi$ turns into a Majorana zero-mode confined to the Josephson vortex-core[26].
+This locality is reflected in the bound-state spectrum, as recently demonstrated in a semiconductor 2DEG-based junction [38].
+If $B$ is strong enough to nucleate Josephson vortices, the position where the local phase difference becomes $\varphi=(2n+1)\pi$ is identified as the Josephson vortex-core, and in TI Josephson junctions, the gapless 1D mode in the Fu-Kane theory for global $\varphi=(2n+1)\pi$ turns into a Majorana zero-mode confined to the Josephson vortex-core [27].
 The concept of this Majorana zero-mode generation is sketched in Fig.
 3a.
-A spatially varying phase $\varphi$ along a TI Josephson junction caused by $B$ leads to a position-dependent gap function $m\propto\cos\frac{\varphi(y)}{2}$ as indicated by the colour code in the junction[26].
+A spatially varying phase $\varphi$ along a junction caused by $B$ leads to a position-dependent gap function $m\propto\cos\frac{\varphi(y)}{2}$ as indicated by the color code [27].
 Associated with the sign change in the gap function, the gap closes locally at $\varphi(y)=(2n+1)\pi$ (white area), corresponding to the core of the Josephson vortex, and a Majorana zero-mode (green dots) nucleates that is confined along the junction by the phase gradient.
 We note that the topologically-protected nature of the gap closing is intuitively understandable in this situation: The region of the junction with $-\pi<\varphi(y)<\pi$ (red) has the even-parity ground state with $\mathbb{Z}_{2}$ index of 0, whereas the region with $\pi<\varphi(y)<3\pi$ (blue) has the odd-parity ground state with $\mathbb{Z}_{2}$ index of 1; therefore, the boundary between the two regions must become gapless for the topological invariant to change.
+Figure 4: Stability of the gap closing and reopening against a change in the chemical potential.
+a, Differential conductance at a high bias voltage of 1 mV as a function of globally-applied back-gate voltage $V_{\rm G}$.
+For both tunnel contacts TJ1 and TJ2, the overall conductance is reduced for more negative gate voltages, indicating an $n$-type regime.
+The vertical dashed lines denote the gate voltages for which the continuous evolution of the spectrum is shown in panels c and d.
+b, Normalized $dI/dV$, denoted $\left[\frac{dI}{dV}\right]_{\rm norm}$, for TJ1 mapped in the $V_{\rm b}$ vs $V_{\rm G}$ plane for $\varphi=0$ (upper panel) and $\varphi=\pi$ (lower panel).
+The normalization was done at each $V_{\rm G}$ by using the $dI/dV$ value at $V_{\rm b}=0.3$ mV as reference.
+c,d, Evolution of the $B$-dependence of the spectrum for a wide range of $V_{\rm G}$ from $-20$ to $+20$ V for TJ1 (c) and TJ2 (d).
+The $V_{\rm G}$ values selected for the presentation are equally spaced.
 We indeed observed the locality of this gap closing.
@@
 This difference in the period is consistent with the sample geometry (see supplement for details) in which the effective junction area considering the flux focusing effect [37, 38, 39] is about $\sim$10of the loop area.
-To understand the LDOS at TJ1 and TJ2, we have calculated the expected texture of the gap function based on the device geometry.
+To understand the LDOS at TJ1 and TJ2, we calculated the expected texture of the gap function based on the device geometry.
 In our junctions, the self-inductance is negligible and its width is much shorter than the Josephson penetration depth (see supplement), justifying the description of the phase gradient in terms of
@@
 This demonstrates the local nature of the gap closing in the presence of Josephson vortices.
-The data shown in Fig.
+Note that in a similar experiment on a non-topological Josephson junction [38], the gap was found to remain open at the Josephson vortex-core.
+Our data shown in Fig.
 3c was taken at an applied back-gate voltage of $V_{\rm G}=-5$ V, and we chose the magnetic field range of 5 – 6 mT for the visibility of the phase gradient.
@@
 3b, we show the magnitude of the gap function near a Josephson vortex-core and the finite-energy bound states calculated from the model in Ref.
-26, along with the spatial extent of the expected Majorana zero-mode.
+27, along with the spatial extent of the expected Majorana zero-mode.
 For the parameters relevant to our device with $\mu$ = 25 meV, the calculated finite-energy bound states form a quasi-continuum with a minigap of only 2% of $\Delta$ (i.e.
 3 $\mu$eV), which is of the order of the thermal energy at the base temperature.
-5 Stability of the gap closing
+V Stability of the gap closing
 In the next step, we check the robustness of the observed features with respect to the chemical potential, as tuned by a global back gate.
 Figure 4a shows the $dI/dV$ data in $B$ = 0 T at a high bias voltage ($V_{\rm b}$ = 1 mV) as a function of the back-gate voltage $V_{G}$ for both TJ1 and TJ2.
-The overall tendency is that the conductance decreases for more negative gate-voltages, indicating that the TI surface is in the $n$-type regime.
+The overall tendency is that the conductance decreases for more negative gate-voltages, indicating that the TSS is in the $n$-type regime.
 Superimposed to this general trend, we observe broad peaks (typical peak width of a few V), which are commonly found in dielectric-based tunnelling devices and conventionally explained by resonant states in the dielectric [31].
-Figure 4b presents the $V_{\rm G}$-dependencies of the normalized conductance spectra obtained at TJ1: the upper and lower panels show the evolutions of the maximum excitation gap for $\varphi=0$ (upper) and the spectrum observed for $\varphi=\pi$ (lower) in a wide range of $V_{\rm G}$ from $-20$ to $+20$ V.
+We can use their dispersion to estimate the relation between $V_{G}$ and $\mu$ (see supplement) and find $\Delta\mu/\Delta V_{G}\approx 1-2\,\mathrm{meV/V}$, consistent with results reported in our earlier work [37].
+This implies that the chemical potential $\mu$ changes by at least $40\,\mathrm{meV}$ in the gating range studied here.
+Figure 4b presents the $V_{\rm G}$-dependencies of the normalized conductance spectra obtained at TJ1.
+Here, we show the evolutions of the maximum excitation gap for $\varphi=0$ (upper panel) and the spectrum observed for $\varphi=\pi$ (lower panel) in a wide range of $V_{\rm G}$ from $-20$ to $+20$ V.
 The maximum excitation gap is very stable, indicating that the resonance states due to impurities are not affecting the proximity-induced gap.
-On the other hand, the position of the conductance peak for $\varphi=\pi$ is slightly shifted to the negative $V_{\rm b}$ side at $V_{\rm G}\approx$ 12 – 20 V, which is probably due to an influence of the strong resonance states located around 13 V (see Fig.
-4a).
-Remarkably, the broad ZBCP that accompanies the gap closing is stable for a wide range of $V_{\rm G}$ from $-10$ V to $+12$ V.
-It is worth noting that the ZBCP observed in this $V_{\rm G}$ range is not correlated with the weaker resonance states seen in Fig.
-4a, which further supports the intrinsic origin of the gap closing.
-For $V_{\rm G}\lesssim-10$ V, the gap closing is not accompanied by a clear ZBCP.
-We believe that this is because our non-tunable tunnel contacts did not couple sufficiently well to the bound states in this regime, preventing us from resolving a peak.
+The gap closing at $\varphi=\pi$ is also very stable and is observed at all $V_{\rm G}$ values.
+Note that the occurrence of the broad ZBCP and the details of the LDOS near zero-bias appear to depend on $V_{\rm G}$, but as we noted in Sec.
+III, these features are expected to be sensitive to the details of the experimental conditions; for example, the $dI/dV$ peak is slightly shifted to the negative $V_{\rm b}$ side at $V_{\rm G}\approx$ 12 – 20 V, which is probably due to an influence of the strong resonance state located around 13 V.
+What is important in these data is the robustness of the gap closing, which is observed for the whole range of $V_{\rm G}$.
 In Figs.
@@
 One can immediately see that the basic magnetic-field-periodic change in $dI/dV$ is very stable at all $V_{\rm G}$ values.
-However, upon close inspection, the occasional breaking of exact electron-hole symmetry is visible.
-Particularly in TJ1, the tunnel-coupling of the electrode to the electron and hole states in the TI surface is modified by back gating: At $V_{\rm G}=-10\,\mathrm{V}$, the electron-like bound states give a stronger signal than the hole-like ones, whereas the situation is reversed for $V_{\rm G}=+15\,\mathrm{V}$.
-Similar behaviour has been reported for comparable devices, e.g., based on carbon nanotubes[30], and it was explained to be due to state-dependent tunnel couplings.
-Overall, we conclude that the gap closing is robust with respect to changes of the chemical potential indicating that we observe a topological phase-transition.
-In passing, we mention that in a device where the tunnel barrier was not in a favourable condition to probe the TI surface, we observed strongly dispersing bound states in a very narrow range of $V_{\rm G}$ associated with a spurious ZBCP occurring at a particular $V_{\rm G}$ value (see supplement).
-These features are consistent with the Andreev quantum-dot states predicted in Ref.
-37, which might occur in the dielectric.
-The observations of two different types of ZBCPs, stable and spurious, give confidence that the former is associated with an intrinsic gap closing.
-6 Discussions
-Our observation of a stable gap closing and reopening in a TI Josephson junction having only a moderate transparency cannot be understood in terms of conventional Andreev bound states and gives the first direct proof of the topological phase-transition predicted by the Fu-Kane theory.
-To firm up the appearance of topological superconductivity, it would be an interesting future experiment to employ a charge sensing technique[38] in order to detect the parity change in the junction before and after the gap closing.
-Of course, the definitive proof of topological superconductivity is the observation of Majorana zero-modes, well-protected by a substantial excitation gap, and the demonstration of their non-Abelian exchange statistics through braiding.
+Even though occasional breaking of exact electron-hole symmetry is visible upon close inspection, similar behavior has been reported for comparable devices, e.g., based on carbon nanotubes [33], and it was explained to be due to state-dependent tunnel couplings.
+Overall, we conclude that the gap closing is robust with respect to changes of the chemical potential, consistent with the TSS origin.
+VI Discussions
+It is prudent to mention that a gap closing at $\varphi=\pi$ can occur in some non-topological scenarios.
+A quantum-dot state, symmetrically coupled to SC electrodes and tuned to resonance, leads to a zero-energy state [40].
+This highly fine-tuned case was in fact observed by us in a device where the tunnel barrier was not in a favourable condition to probe the TSS.
+We found strongly dispersing bound states in a very narrow range of $V_{\rm G}$ associated with a spurious ZBCP occurring at a particular $V_{\rm G}$ value (see supplement).
+These features are consistent with the Andreev quantum-dot states [40] which might occur in the dielectric.
+The robust gap closing discussed in the previous section is clearly of different origin.
+Nonetheless, the sharp Andreev quantum-dot states are useful for estimating the energy resolution of our setup to be about 30 $\mu$eV (see supplement).
+The spurious ZBCP in this case also tells us that one should not look for a ZBCP but should focus on the gap closing to elucidate the robustness of the phenomenon.
+Figure 5: Stability of the Majorana zero-modes in the putative case with Nb electrodes.
+a,b, Calculated dependencies of the Majorana bound state width $\lambda_{\rm B}$ and the first excited state energy $E_{1}$ on the chemical potential $\mu$ and the perpendicular magnetic field $B$ for the Majorana zero-mode generated at the Josephson vortex-core in a putative device with Nb as the SC electrodes, which will sustain high magnetic fields.
+We used $L$ = 50 nm, $v_{\rm F}$ = 5.5 $\times$ 105 m/s, and $\Delta$ = 1 meV.
+The blue and green lines in panel b are the boundaries to guarantee a sufficient separation of Majorana zero-modes ($\lambda_{\rm B}/l_{B}<0.2$ obtained from panel a) and a sufficiently large excitation gap ($E_{1}/\Delta>0.2$), respectively.
+Majorana zero-modes are well-protected in the relatively broad region between the blue and green lines.
+In the literature, the gap closing at $\varphi=\pi$ has been reported for diffusive long SNS junctions [41, 32].
+In such a system, disorder localizes most channels, but it also leads to the appearance of a few high-transmission channels [42], which can be responsible for a gap closing.
+In the case of our junctions, the mean-free path $l_{e}\approx$ 40 nm has been estimated for bulk-insulating Bi1.5Sb0.5Te1.7Se1.3 crystals [43], which is a bit shorter but comparable to our junction length (50 nm), putting our device in a quasi-ballistic regime.
+(Remember that our device is in the short-junction limit where the Fu-Kane theory is applicable.) It is possible that some nearly-longitudinal Andreev modes with large $k_{y}$ traveling for a long distance present the gap closing phenomenology of the diffusive scenario.
+Nevertheless, such a gap closing occurs on top of the intrinsic gap closing guaranteed by the spin-momentum locking for the transverse mode [26].
+Hence, the consideration of the diffusive contribution does not change the conclusion that the gap closing predicted by Fu and Kane is being observed in our device, unless the junction is not coupled to the TSS (which can be dismissed based on Ref.
+37).
+To better address the question of which states contribute to what extent to the gap-closing, employing high-frequency measurements to achieve a higher resolution [44] would be useful.
+Furthermore, employing a charge-sensing technique [45] to detect the parity change in the junction before and after the gap closing would unambiguously prove the topological phase transition.
+A definitive proof of Majorana zero-modes is the demonstration of their non-Abelian exchange statistics through braiding.
 In a TI Josephson junction, well-protected Majorana zero-modes can in principle be generated in Josephson vortex-cores as we discuss below.
@@
 For large $\mu\gg\Delta$, the formula goes over to the simple expression $v_{\rm M}\simeq v_{\rm F}(\Delta/\mu)^{2}$ given in Ref.
-26.
+27.
 The other important quantity is the energy of the first excited state, which is given as $E_{1}\approx\sqrt{2\pi v_{\rm M}\Delta/l_{B}}$ near $\varphi=(2n+1)\pi$ [27].
@@
 The results, shown in Figs.
-2c and 2d, indicate that a moderately large $\mu$ is indeed helpful to realize a short $\lambda_{\rm B}$ and the magnetic field helps to obtain a usefully large $E_{1}$.
+5a and 5b, indicate that a moderately large $\mu$ is indeed helpful to realize a short $\lambda_{\rm B}$ and the magnetic field helps to obtain a usefully large $E_{1}$.
 In Fig.
-2d, we draw the boundaries to guarantee a sufficient separation of Majorana zero-modes ($\lambda_{\rm B}/l_{B}<0.2$, blue line) and a sufficiently large excitation gap ($E_{1}/\Delta>0.2$, green line).
+5b, we draw the boundaries to guarantee a sufficient separation of Majorana zero-modes ($\lambda_{\rm B}/l_{B}<0.2$, blue line) and a sufficiently large excitation gap ($E_{1}/\Delta>0.2$, green line).
 In the region between the blue and green lines, see Fig.
-2d, Majorana zero-modes are well-protected.
-In fact, these estimates suggest that it is not necessary to tune $\mu$ very close to the Dirac point as long as a magnetic field of the order of 100 mT is used to generate Josephson vortices — an encouraging result for the prospects of Majorana zero-modes in this platform.
+5b, Majorana zero-modes are well-protected.
+In fact, these estimates suggest that it is not necessary to tune $\mu$ very close to the Dirac point as long as a magnetic field of the order of 100 mT is used to generate Josephson vortices — an encouraging result for this platform.
 It is prudent to note that the soft gap, which we observed in our devices, is a problem for braiding, and a hard gap must be realized by improving the interface transparency between the TI and a superconductor.
-In this regard, we envision the use of the self-epitaxy technique discovered for proximitising (Bi1-xSbx)2Te3 thin films with the PdTe2 superconductor, because a very high transparency of 0.96 was achieved with this technique[41].
-In conclusion, we have observed periodic gap-closing and reopening in a phase-biased TI Josephson junction via tunnelling spectroscopy.
-We have shown that this phenomenon is robust as a function of an extended parameter range in the chemical potential and the perpendicular magnetic field.
-The robustness of the observed gap closing and reopening is in sharp contrast to the signal expected for a conventional moderate-transparency Josephson junction or other types of zero-energy states which originate from Andreev quantum-dot states[37].
-The robustness of the gap closing and reopening is a key feature of the topological phase-transition that occurs in the 1D Majorana state as a function of $\varphi$.
-In the magnetic-field range where a couple of Josephson vortices nucleate in the junction, we have observed the locality of the gap closing implying a confinement of the zero-energy mode[26].
-However, within the magnetic-field range where Al remains superconducting, an abundance of low-lying excited states arises and this leads to spectra that are indistinguishable from the gap closing of an extended 1D Majorana mode.
+In conclusion, we have observed periodic gap-closings and reopenings in a phase-biased TI Josephson junction, which is consistent with the existence of perfectly-transmitted Andreev modes guaranteed by the spin-momentum locking of the TSS.
+This phenomenon is robust as a function of an extended parameter range in the chemical potential and the perpendicular magnetic field, which support its TSS origin.
+The spin-non-degeneracy of the TSS makes this gap-closing and reopening to be topological phase transition accompanied by parity change.
+In the magnetic-field range where a couple of Josephson vortices nucleate in the junction, we observed the locality of the gap closing implying a confinement of the zero-energy mode [27].
+However, within the magnetic-field range where Al remains superconducting, an abundance of low-lying excited states arises and this leads to spectra that are indistinguishable from the gap closing near zero field.
 We have discussed the necessary parameter range to realize well-protected Majorana zero-modes in the Josephson vortex core.
-Intriguingly, we found that maintaining a chemical potential $\mu$ not too close to the Dirac point is crucial to avoid Majorana delocalisation; a $\mu$ of approximately 25 meV facilitates the realization of well-protected Majorana zero-modes in a magnetic field of 100 mT accessible with a Nb electrode.
+Intriguingly, we found that maintaining a chemical potential $\mu$ not too close to the Dirac point is crucial for avoiding Majorana delocalisation; a $\mu$ of approximately 25 meV facilitates the realization of well-protected Majorana zero-modes in a magnetic field of 100 mT accessible with a Nb electrode.
 Considering the high controllability of the Josephson vortex, this result is promising for braiding experiments.
-Overall, our experiment elucidates the topological nature of a TI Josephson junction and underscores the realistic prospect of generating well-protected Majorana zero-modes in this system.
-{methods}
-Materials and device fabrications: Bulk single crystals of BiSbTeSe2 were synthesized by the modified Bridgman method using high-purity (99.9999%) Bi, Sb, Te, and Se as starting materials by following the recipe described in Ref.
-28.
-We mechanically exfoliated ultra-thin BiSbTeSe2 flakes from bulk single crystals on degenerately-doped Si wafers covered by 280 nm SiO2.
-The latter acts as dielectric for the back gate.
-BiSbTeSe2 flakes suitable for device fabrication were identified under an optical microscope.
-The junctions were defined by electron beam lithography.
-PMMA A4 resist was exposed using a Raith PIONEER Two system.
-After development, residual resist in the contact areas was removed by a gentle oxygen plasma treatment.
-Subsequent cleaning in a hydrochloric acid solution shortly before metallisation ensured the removal of any oxide layer.
-Then, 3 nm + 45 nm Ti/Al contacts were thermally evaporated after a gentle in-situ Ar cleaning and the device was exposed to air to form a native Al2O3 layer.
-An additional atomic layer deposition of a thin Al2O3 layer onto the whole device was performed at 80 C using Ultratec Savannah S200 to form a tunnel barrier.
-For the fabrication of the tunnel electrodes, another electron beam lithography was performed to define the pattern.
-A 5-nm-thick Pt layer was sputter-deposited to act as wetting layer, followed by sputtering an additional 45-nm-thick Au layer.
-The contact resistances achieved by this procedure are typically larger than $100\mathrm{k\Omega}$, ensuring electron tunnelling to be the prevalent form of transport.
-The precise device geometry was determined by scanning electron microscopy after the measurements were finished.
-Measurements: Transport measurements were performed at base temperature ($\sim$50 mK) of our dry dilution refrigerator (Oxford Instruments TRITON 300).
-Electrical lines incorporate RC and copper-powder filters to reduce noise.
-Using a two-probe configuration, the differential conductance $dI/dV$ was measured between each individual tunnel electrode and the SC electrode with a standard low-frequency lock-in technique using a Basel instruments current preamplifier.
-A small DC-offset introduced by the preamplifier was numerically corrected for after the measurements.
-An AC voltage of 5–10 $\mu$V was superimposed to a DC bias voltage.
-For investigations of the magnetic field dependence, we used our integrated 6/1/1-T superconducting vector magnet.
-Simulations: To qualitatively reproduce the conductance spectrum of the TI Josephson junction, a transport simulation was performed using the Python library Kwant[42].
-As a model for the three-dimensional strong TI, the Bernevig-Hughes-Zhang model was chosen and implemented on a cubic lattice.
-Superconductivity was included in the Bogoliubov–de-Gennes formalism.
-The system has been extended infinitely along the junction and probed by a tunnel lead attached perpendicularly to the junction.
-For more information see supplement.
-{addendum}
+Overall, our experiment demonstrates a highly promising nature of the TI Josephson junctions for the Majorana research thanks to the naturally-existing spin-momentum locking of the TSS.
+Data availability: The data and codes that support the findings of this study are available at the online depository zenodo with the identifier 10.5281/zenodo.11615512 and Supplemental Material.
+Acknowledgements.
+We thank Liang Fu, Max Geier, Michal Papaj, and Henry Legg for fruitful discussions.
 This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant Agreement No.
 741121) and was also funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1 - 390534769, as well as by the DFG under CRC 1238 - 277146847 (Subprojects A04 and B01).
-MB grew the BSTS2 crystals, JS and ENN fabricated the devices, performed the experiments and analysed the data together with YA.
-AM and AZ performed numerical calculations, FH and YA conceived and supervised the project.
-JS, FH, and YA wrote the manuscript with input from all authors.
-The authors declare no competing interests.
-Correspondence and requests for materials should be addressed to Y.A. (ando@ph2.uni-koeln.de).
-Data availability: The data that support the findings of this study are available at the online depository zenodo with the identifier 10.5281/zenodo.11615514 and in Supplementary Information.